Two cars are on a race track. It takes car "A" 1.5 minutes to drive a lap. It takes car "B" 2 minutes to drive a lap. How many minutes does it take for car A to pass car B?
Answer by checkley71(8403)

1/1.5(X+1)=1/2X
1/(1.5X+1.5)=1/2X
2X=1.5X+1.5
2X-1.5X=1.5
.5X=1.5
X=1.5/.5
X=3 LAPS FOR THE "B" CAR
PROOF
3*2=(3+1)1.5
6=4*1.5
6=6 MINUTES WHEN THE PASS